# Large Language Models: A Survey

### v2

@@ I Introduction
The recent advances on transformer-based large language models (LLMs), pretrained on Web-scale text corpora, significantly extended the capabilities of language models (LLMs). For example, OpenAI’s ChatGPT and GPT-4 can be used not only for natural language processing, but also as general task solvers to power Microsoft’s Co-Pilot systems, for instance, can follow human instructions of complex new tasks performing multi-step reasoning when needed. LLMs are thus becoming the basic building block for the development of general-purpose AI agents or artificial general intelligence (AGI).
@@
Large language models (LLMs) mainly refer to transformer-based neural language models  [^1] Recently, several very promising non-transformer LLMs have been proposed, such as the LLMs based on structured state space models [29, 30]. See Section VII for more details. that contain tens to hundreds of billions of parameters, which are pre-trained on massive text data, such as PaLM [31], LLaMA [32], and GPT-4 [33], as summarized in Table III. Compared to PLMs, LLMs are not only much larger in model size, but also exhibit stronger language understanding and generation abilities, and more importantly, emergent abilities that are not present in smaller-scale language models. As illustrated in Fig. 1, these emergent abilities include (1) in-context learning, where LLMs learn a new task from a small set of examples presented in the prompt at inference time, (2) instruction following, where LLMs, after instruction tuning, can follow the instructions for new types of tasks without using explicit examples, and (3) multi-step reasoning, where LLMs can solve a complex task by breaking down that task into intermediate reasoning steps as demonstrated in the chain-of-thought prompt [34]. LLMs can also be augmented by using external knowledge and tools [35, 36] so that they can effectively interact with users and environment [37], and continually improve itself using feedback data collected through interactions (e.g. via reinforcement learning with human feedback (RLHF)).
@@ II Large Language Models
BERT
110M, 340M
2018
137B
BooksCorpus, English Wikipedia
RoBERTa
355M
2019
2.2T
BooksCorpus, English Wikipedia, CC-NEWS, STORIES (a subset of Common Crawl), Reddit
Encoder-Only
ALBERT
12M, 18M, 60M, 235M
2019
137B
BooksCorpus, English Wikipedia
DeBERTa
2020
BooksCorpus, English Wikipedia, STORIES, Reddit content
XLNet
110M, 340M
2019
32.89B
BooksCorpus, English Wikipedia, Giga5, Common Crawl, ClueWeb 2012-B
Decoder-only
GPT-1
120M
2018
1.3B
BooksCorpus
GPT-2
1.5B
2019
10B
Reddit outbound
T5 (Base)
223M
2019
156B
Common Crawl
Encoder-Decoder
MT5 (Base)
300M
2020
New Common Crawl-based dataset in 101 languages (m Common Crawl)
BART (Base)
139M
2019
Corrupting text
GPT-3
125M, 350M, 760M, 1.3B, 2.7B, 6.7B, 13B, 175B
2020
$\times$
300B
Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia
GPT Family
CODEX
12B
2021
GPT
Public GitHub software repositories
WebGPT
760M, 13B, 175B
2021
GPT-3
$\times$
ELI5
GPT-4
1.76T
2023
$\times$
13T
LLaMA1
7B, 13B, 33B, 65B
2023
1T, 1.4T
Online sources
LLaMA2
7B, 13B, 34B, 70B
2023
2T
Online sources
Alpaca
7B
2023
LLaMA1
GPT-3.5
Vicuna-13B
13B
2023
LLaMA1
GPT-3.5
LLaMA Family
Koala
13B
2023
LLaMA
Dialogue data
Mistral-7B
7.3B
2023
Code Llama
34
2023
LLaMA2
500B
Publicly available code
LongLLaMA
3B, 7B
2023
OpenLLaMA
1T
LLaMA-Pro-8B
8.3B
2024
LLaMA2-7B
80B
Code and math corpora
TinyLlama-1.1B
1.1B
2024
LLaMA1.1B
3T
SlimPajama, Starcoderdata
PaLM
8B, 62B, 540B
2022
$\times$
780B
Web documents, books, Wikipedia, conversations, GitHub code
U-PaLM
8B, 62B, 540B
2022
$\times$
1.3B
Web documents, books, Wikipedia, conversations, GitHub code
PaLM Family
PaLM-2
340B
2023
3.6T
Web documents, books, code, mathematics, conversational data
Med-PaLM
540B
2022
PaLM
$\times$
780B
HealthSearchQA, MedicationQA, LiveQA
Med-PaLM 2
2023
PaLM 2
$\times$
MedQA, MedMCQA, HealthSearchQA, LiveQA, MedicationQA
FLAN
137B
2021
LaMDA-PT
Web documents, code, dialog data, Wikipedia
Gopher
280B
2021
$\times$
300B
MassiveText
ERNIE 4.0
10B
2023
$\times$
4TB
Chinese text
Retro
7.5B
2021
$\times$
600B
MassiveText
LaMDA
137B
2022
$\times$
168B
public dialog data and web documents
ChinChilla
70B
2022
$\times$
1.4T
MassiveText
Galactia-120B
120B
2022
450B
Other Popular LLMs
CodeGen
16.1B
2022
THE PILE, BIGQUERY, BIGPYTHON
BLOOM
176B
2022
366B
ROOTS
Zephyr
7.24B
2023
Mistral-7B
800B
Synthetic data
Grok-0
33B
2023
$\times$
Online source
ORCA-2
13B
2023
LLaMA2
2001B
StartCoder
15.5B
2023
35B
GitHub
MPT
7B
2023
1T
RedPajama, m Common Crawl, S2ORC, Common Crawl
Mixtral-8x7B
46.7B
2023
Instruction dataset
Falcon 180B
180B
2023
3.5T
RefinedWeb
Gemini
1.8B, 3.25B
2023
Web documents, books, and code, image data, audio data, video data
DeepSeek-Coder
1.3B, 6.7B, 33B
2024
2T
GitHub’s Markdown and StackExchange
DocLLM
1B,7B
2024
$\times$
2T
IIT-CDIP Test Collection 1.0, DocBank
@@ II-B1 The GPT Family
GPT-3 [56] is a pre-trained autoregressive language model with 175 billion parameters. GPT-3 is widely considered as the first LLM in that it not only is much larger than previous PLMs, but also for the first time demonstrates emergent abilities that are not observed in previous smaller PLMs. GPT-3 shows the emergent ability of in-context learning, which means GPT-3 can be applied to any downstream tasks without any gradient updates or fine-tuning, with tasks and few-shot demonstrations specified purely via text interaction with the model. GPT-3 achieved strong performance on many NLP tasks, including translation, question-answering, and the cloze tasks, as well as several ones that require on-the-fly reasoning or domain adaptation, such as unscrambling words, using a novel word in a sentence, 3-digit arithmetic. Fig 9 plots the performance of GPT-3 as a function of the number of examples in in-context prompts.
@@ III-I2 Low-Rank Adaption (LoRA)
It is worth mentioning that LoRA can be applied to any a subset of weight matrices in a neural network to reduce the number of trainable parameters. In the Transformer architecture, there are four weight matrices in the self-attention module ($W_{q}$ , $W_{k}$, $W_{v}$ , $W_{o}$), and two in the MLP module. Most of the time, LoRA is focused on adapting the attention weights only for downstream tasks, and freezes the MLP modules, so they are not trained in downstream tasks both for simplicity and parameter-efficiency.
@@ IV-A LLM limitations
  • They don’t have state/memory. LLMs on their own cannot remember even what was sent to them in the previous prompt. That is an important limitation for many of the uses cases that require some form of state.
@@ V-C Datasets for Augmented: using external knowledge/tools
HumanEval
PASS@k
Link
Link
Link
MBPP
PASS@k, Accuracy
Link
Link
APPS
PASS@k, Accuracy
Link
Link
WikiSQL
Accuracy
Link
Link
CoNaLa
BLEU
Link
Link
CodeParrot
PASS@k
Link
HellaSwag
Accuracy
Link
Link
Link
AI2 Reasoning Challenge (ARC)
Accuracy
Link
Link
Link
BoolQ
Accuracy
Link
Link
MultiRC
F1-score, Accuracy
Link
Link
CNN/Daily Mail [200]
Accuracy
Link
SQuAD
F1-score, EM
Link
Link
Link
RACE
Accuracy
Link
Link
CNN/Daily Mail [201]
ROUGE
Link
Link
Drop
F1-score, EM
Link
Link
Link
QuAC
F1-score, HEQ-Q, HEQ-D
Link
Link
Link
TriviaQA
EM, F1-score, Accuracy
Link
Link
Link
Natural Questions
EM, F1-score, Accuracy
Link
Link
Link
StrategyQA
Accuracy, Recall@10, SARI
Link
Link
Link
CoQA
F1-score
Link
Link
Link
XSum
ROUGE
Link
Link
SAMSum
ROUGE
Link
WikiSum
ROUGE
Link
DialogSum
ROUGE
Link
Link
TruthfulQA
MC1 , MC2, % true, % info, BLEURT
Link
Link
Link
MMLU
Accuracy
Link
Link
Link
GSM8K
Accuracy
Link
Link
Link
PIQA
Accuracy
Link
Link
Link
SIQA
Accuracy
Link
Link
Link
OpenBookQA (OBQA)
Accuracy
Link
Link
Link
HotpotQA
EM, F1-score, Joint EM, Joint F1-score,
Link
Link
Link
MATH
Accuracy
Link
Link
CommonsenseQA
Accuracy
Link
Link
Link
Natural Instructions
ROUGE-L, Human
Link
Link
Link
BIG-bench
Accuracy, Average
Link
Link
ToolTalk
Success rate, Precision, Recall, Incorrect action rate, Percent of failing error types
Link
Link
MetaTool
Accuracy, Precision, Recall, F1-score
Link
Link
GPT4Tools
Successful Rate of Thought, Successful Rate of Action, Successful Rate of Arguments, Success Rate
Link
Link
API-Bank
Correctness, ROUGE, Error(API Hallucination, Has Exception, Invalid Input Parameters, False API Call Format, API Call, Miss Input Parameters)
Link
Link
Alpaca-CoT
Link
Link
@@ VII-E Security and Ethical/Responsible AI
Ensuring the robustness and security of LLMs against adversarial attacks and other vulnerabilities is a critical area of research [219]. As LLMs are increasingly deployed in real-world applications, they need to be protected from potential threats, to prevent them being used to manipulate people or spread mis-information.

### v3

@@ I Introduction
The recent advances on transformer-based large language models (LMs), pretrained on Web-scale text corpora, significantly extended the capabilities of language models (LLMs). For example, OpenAI’s ChatGPT and GPT-4 can be used not only for natural language processing, but also as general task solvers to power Microsoft’s Co-Pilot systems, for instance, can follow human instructions of complex new tasks performing multi-step reasoning when needed. LLMs are thus becoming the basic building block for the development of general-purpose AI agents or artificial general intelligence (AGI).
@@
Large language models mainly refer to transformer-based neural language models  [^1] Recently, several very promising non-transformer LLMs have been proposed, such as the LLMs based on structured state space models [29, 30]. See Section VII for more details. that contain tens to hundreds of billions of parameters, which are pre-trained on massive text data, such as PaLM [31], LLaMA [32], and GPT-4 [33], as summarized in Table III. Compared to PLMs, LLMs are not only much larger in model size, but also exhibit stronger language understanding and generation abilities, and more importantly, emergent abilities that are not present in smaller-scale language models. As illustrated in Fig. 1, these emergent abilities include (1) in-context learning, where LLMs learn a new task from a small set of examples presented in the prompt at inference time, (2) instruction following, where LLMs, after instruction tuning, can follow the instructions for new types of tasks without using explicit examples, and (3) multi-step reasoning, where LLMs can solve a complex task by breaking down that task into intermediate reasoning steps as demonstrated in the chain-of-thought prompt [34]. LLMs can also be augmented by using external knowledge and tools [35, 36] so that they can effectively interact with users and environment [37], and continually improve itself using feedback data collected through interactions (e.g. via reinforcement learning with human feedback (RLHF)).
@@ II Large Language Models
BERT 110M, 340M 2018 - 137B BooksCorpus, English Wikipedia
RoBERTa 355M 2019 - 2.2T BooksCorpus, English Wikipedia, CC-NEWS, STORIES (a subset of Common Crawl), Reddit
Encoder-Only ALBERT 12M, 18M, 60M, 235M 2019 - 137B BooksCorpus, English Wikipedia
DeBERTa - 2020 - - BooksCorpus, English Wikipedia, STORIES, Reddit content
XLNet 110M, 340M 2019 - 32.89B BooksCorpus, English Wikipedia, Giga5, Common Crawl, ClueWeb 2012-B
Decoder-only GPT-1 120M 2018 - 1.3B BooksCorpus
GPT-2 1.5B 2019 - 10B Reddit outbound
T5 (Base) 223M 2019 - 156B Common Crawl
Encoder-Decoder MT5 (Base) 300M 2020 - - New Common Crawl-based dataset in 101 languages (m Common Crawl)
BART (Base) 139M 2019 - - Corrupting text
GPT-3 125M, 350M, 760M, 1.3B, 2.7B, 6.7B, 13B, 175B 2020 $\times$ 300B Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia
GPT Family CODEX 12B 2021 GPT - Public GitHub software repositories
WebGPT 760M, 13B, 175B 2021 GPT-3 $\times$ - ELI5
GPT-4 1.76T 2023 - $\times$ 13T -
LLaMA1 7B, 13B, 33B, 65B 2023 - 1T, 1.4T Online sources
LLaMA2 7B, 13B, 34B, 70B 2023 - 2T Online sources
Alpaca 7B 2023 LLaMA1 - GPT-3.5
Vicuna-13B 13B 2023 LLaMA1 - GPT-3.5
LLaMA Family Koala 13B 2023 LLaMA - Dialogue data
Mistral-7B 7.3B 2023 - -
Code Llama 34 2023 LLaMA2 500B Publicly available code
LongLLaMA 3B, 7B 2023 OpenLLaMA 1T -
LLaMA-Pro-8B 8.3B 2024 LLaMA2-7B 80B Code and math corpora
TinyLlama-1.1B 1.1B 2024 LLaMA1.1B 3T SlimPajama, Starcoderdata
PaLM 8B, 62B, 540B 2022 - $\times$ 780B Web documents, books, Wikipedia, conversations, GitHub code
U-PaLM 8B, 62B, 540B 2022 - $\times$ 1.3B Web documents, books, Wikipedia, conversations, GitHub code
PaLM Family PaLM-2 340B 2023 - 3.6T Web documents, books, code, mathematics, conversational data
Med-PaLM 540B 2022 PaLM $\times$ 780B HealthSearchQA, MedicationQA, LiveQA
Med-PaLM 2 - 2023 PaLM 2 $\times$ - MedQA, MedMCQA, HealthSearchQA, LiveQA, MedicationQA
FLAN 137B 2021 LaMDA-PT - Web documents, code, dialog data, Wikipedia
Gopher 280B 2021 - $\times$ 300B MassiveText
ERNIE 4.0 10B 2023 - $\times$ 4TB Chinese text
Retro 7.5B 2021 - $\times$ 600B MassiveText
LaMDA 137B 2022 - $\times$ 168B public dialog data and web documents
ChinChilla 70B 2022 - $\times$ 1.4T MassiveText
Galactia-120B 120B 2022 - 450B
Other Popular LLMs CodeGen 16.1B 2022 - - THE PILE, BIGQUERY, BIGPYTHON
BLOOM 176B 2022 - 366B ROOTS
Zephyr 7.24B 2023 Mistral-7B 800B Synthetic data
Grok-0 33B 2023 - $\times$ - Online source
ORCA-2 13B 2023 LLaMA2 - 2001B -
StartCoder 15.5B 2023 - 35B GitHub
MPT 7B 2023 - 1T RedPajama, m Common Crawl, S2ORC, Common Crawl
Mixtral-8x7B 46.7B 2023 - - Instruction dataset
Falcon 180B 180B 2023 - 3.5T RefinedWeb
Gemini 1.8B, 3.25B 2023 - Web documents, books, and code, image data, audio data, video data
DeepSeek-Coder 1.3B, 6.7B, 33B 2024 - 2T GitHub’s Markdown and StackExchange
DocLLM 1B,7B 2024 - $\times$ 2T IIT-CDIP Test Collection 1.0, DocBank
@@ II-B1 The GPT Family
GPT-3 [56] is a pre-trained autoregressive language model with 175 billion parameters. GPT-3 is widely considered as the first LLM in that not only it is much larger than previous PLMs, but also for the first time demonstrates emergent abilities that are not observed in previous smaller PLMs. GPT-3 shows the emergent ability of in-context learning, which means GPT-3 can be applied to any downstream tasks without any gradient updates or fine-tuning, with tasks and few-shot demonstrations specified purely via text interaction with the model. GPT-3 achieved strong performance on many NLP tasks, including translation, question-answering, and the cloze tasks, as well as several ones that require on-the-fly reasoning or domain adaptation, such as unscrambling words, using a novel word in a sentence, 3-digit arithmetic. Fig 9 plots the performance of GPT-3 as a function of the number of examples in in-context prompts.
@@ III-I2 Low-Rank Adaption (LoRA)
It is worth mentioning that LoRA can be applied to a subset of weight matrices in a neural network to reduce the number of trainable parameters. In the Transformer architecture, there are four weight matrices in the self-attention module ($W_{q}$ , $W_{k}$, $W_{v}$ , $W_{o}$), and two in the MLP module. Most of the time, LoRA is focused on adapting the attention weights only for downstream tasks, and freezes the MLP modules, so they are not trained in downstream tasks both for simplicity and parameter-efficiency.
@@ IV-A LLM limitations
  • They don’t have state/memory. LLMs on their own cannot remember even what was sent to them in the previous prompt. That is an important limitation for many of the use cases that require some form of state.
@@ V-C Datasets for Augmented: using external knowledge/tools
HumanEval PASS@k Link Link Link
MBPP PASS@k, Accuracy - Link Link
APPS PASS@k, Accuracy - Link Link
WikiSQL Accuracy - Link Link
CoNaLa BLEU Link Link
CodeParrot PASS@k - Link -
HellaSwag Accuracy Link Link Link
AI2 Reasoning Challenge (ARC) Accuracy Link Link Link
BoolQ Accuracy - Link Link
MultiRC F1-score, Accuracy - Link Link
CNN/Daily Mail [200] Accuracy - Link -
SQuAD F1-score, EM Link Link Link
RACE Accuracy - Link Link
CNN/Daily Mail [201] ROUGE - Link Link
Drop F1-score, EM Link Link Link
QuAC F1-score, HEQ-Q, HEQ-D Link Link Link
TriviaQA EM, F1-score, Accuracy Link Link Link
Natural Questions EM, F1-score, Accuracy Link Link Link
StrategyQA Accuracy, Recall@10, SARI Link Link Link
CoQA F1-score Link Link Link
XSum ROUGE - Link Link
SAMSum ROUGE - - Link
WikiSum ROUGE - Link -
DialogSum ROUGE - Link Link
TruthfulQA MC1 , MC2, % true, % info, BLEURT Link Link Link
MMLU Accuracy Link Link Link
GSM8K Accuracy Link Link Link
PIQA Accuracy Link Link Link
SIQA Accuracy Link Link Link
OpenBookQA (OBQA) Accuracy Link Link Link
HotpotQA EM, F1-score, Joint EM, Joint F1-score, Link Link Link
MATH Accuracy - Link Link
CommonsenseQA Accuracy Link Link Link
Natural Instructions ROUGE-L, Human Link Link Link
BIG-bench Accuracy, Average - Link Link
ToolTalk Success rate, Precision, Recall, Incorrect action rate, Percent of failing error types - Link Link
MetaTool Accuracy, Precision, Recall, F1-score - Link Link
GPT4Tools Successful Rate of Thought, Successful Rate of Action, Successful Rate of Arguments, Success Rate - Link Link
API-Bank Correctness, ROUGE, Error(API Hallucination, Has Exception, Invalid Input Parameters, False API Call Format, API Call, Miss Input Parameters) - Link Link
Alpaca-CoT - - Link Link
@@ VII-E Security and Ethical/Responsible AI
Ensuring the robustness and security of LLMs against adversarial attacks and other vulnerabilities is a critical area of research [219]. As LLMs are increasingly deployed in real-world applications, they need to be protected from potential threats, to prevent them being used to manipulate people or spread mis-information. Improving the reasoning capabilities of these model [220], would help them to better detect potential adversarial attacks.